# Vec2Face: Scaling Face Dataset Generation with Loosely Constrained Vectors

### v1

@@ 4.4 Further analysis
Vec2Face generates datasets with large inter-class separability. Fig. 5 presents the number of identities whose cosine similarity against any other identity vectors is less than 0.4, plotted against the number of generated identities in different datasets. A reasonable threshold varies for using different FR models (Kim et al., 2023; Boutros et al., 2024). We select 0.4 is because it gives us a reasonable number (183k out of 200k) on a real dataset. We sample 200K vectors to generate identity images by Vec2Face and Arc2Face, where we use the 4-step sampler (Luo et al., 2023) provided in the Arc2Face repository. We use ArcFace-R100 to extract identity features from WebFace4M (Zhu et al., 2023) and the synthesized identity images. We have two observations.
Figure 5: Inter-class separability measurement. The number of well-separated identities under different identity scales in the synthetic datasets with 10k identities, 200k generated identities from Arc2Face and Vec2Face, and real identities from WebFace4M.
First, synthetic datasets such as SynFace and SFace generally have lower inter-class separability than real-world datasets. This is because their data synthesis methods do not utilize the face feature characteristics for image generation, as Vec2Face and Idiff-Face do. Second, Arc2Face has much larger identity separability than existing methods, comparing with the number reported in the (Kim et al., 2023), but is still inferior to Vec2Face. This is because the initial noise image sometimes exerts a stronger influence on the final output than the conditioned identity vector. Third, real-world datasets have slightly lower identity separability than our method, because they have a small number of twins, close relatives, or even the same person with different character names in TV shows (Wu et al., 2024b). Note that 0.4 is not used to evaluate the number of actual identities in the dataset, but rather to provide a consistent measurement of identity separability across datasets. In fact, the special design of Vec2Face ensures a high rate of transferring the similarity of sampled vectors to the similarity of generated images.
Impact of inter-class separability on accuracy. Table 7 summarizes how different average identity similarity or different levels of identity separability affect FR accuracy. Results indicate that higher identity separability is beneficial, which is consistent with previous finding (Boutros et al., 2023a). Moreover, we find that using a more strict $\tau=0.2$ does reduce a large number of sampled vectors, but it does not further benefit the performance.
@@
Table 5: Impact of $\sigma$ values with corresponding sampling % for each identity on average FR accuracy (%). We use 0.5M images from 10K identities for training. AttrOP is not applied and Simmin is the minimum cosine similarity between the perturbed vectors and their identity vectors.
@@
Table 6: Evaluation on other tasks. Hadrian, Eclipse, SLLFW, and DoppelVer emphasize facial hair variation, face exposure difference, similar-looking, and doppelganger tasks.
@@
Table 7: Impact of inter-class separability. Larger average identity similarity (low separability) reduces the accuracy. We use 0.5M images from 10k identities for training.
@@
Table 8: Impact of number of images per identity. Different number of images are generated by the best $\sigma$ and their sampling % setting. Average accuracy (%) is reported.
Architecture Avg.
@@
Table 9: HSFace10k training on ResNets and Vision Transformers. The average accuracy (%) is reported.
Vec2Face generates datasets with large intra-class variation. Table 1 indicates the generated dataset has large variation on age and pose. We also test the performance on the test sets emphasized on the variation of other attributes. Hadrian and Eclipse (Wu et al., 2024a) have 6,000 image pairs for each and respectively control the variation on facial hair and face exposure. Table 6 shows that, while HSFace10k underperforms on accuracy, increasing the dataset size eventually surpasses the accuracy of the real dataset.
Impact of hyperparameter $\sigma$. In Vec2Face, $\sigma$ directly controls intra-class variation, i.e., larger $\sigma$ means higher intra-class variation and vice versa. We test different configurations of $\sigma$ values and its sampling %. Table 5 shows that, When increasing $\sigma$, FR accuracy first increases and then drops. Initially, increasing intra-class variation is beneficial because it creates some hard training samples. But if $\sigma$ is too large, the generated face may look like a different person, degrading performance. Same for increasing the sampling % of the large $\sigma$. Thus we choose $\sigma=\{0.3,0.5,0.7\}$ to sample $\{40\%,40\%,20\%\}$ images for each identity.
Does Vec2Face generate new identities compared with its training identities? We compute the feature similarity between the 300k identities from the proposed synthetic data and the 50k training identities selected from WebFace4M. We find that only 0.006%, or 168 out of the 300K synthetic identities are similar to the training identities, with a similarity score greater than 0.4. To completely avoid using real identities for training FR models, we have removed these 168 identities from HSFace300K, and the performance change is negligible.
More images per identity is beneficial. In Table 8, increasing the number of images per identity from 5 to 80 increases the accuracy. The improvement is gradually saturated.
Larger backbone helps the model performance. In Table 9, accuracy of four models trained with HSFace10K and ArcFace loss increases from SE-IResNet18 to SE-IResNet100 and is saturated at SE-IResNet200. It is also useful for training with Vision Transformer (Dosovitskiy, 2020) backbones.
Limitations. The proposed synthetic dataset is effective for general FR but would be much less useful for fine-grained tasks such as discriminating between doppelgangers. See Table 6. We believe a better understanding of ‘identity’ and stronger generation models will help solve this issue and leave this to future work.
@@ 5 Conclusions
This paper proposes Vec2Face, a face generation method which can elegantly create a large number of identities and face images. This method conveniently converts vectors to face images and can translate perturbations of the vectors into consistent perturbations of the resulting face images. Because of its unique design, the inter-class and intra-class variations of generated face images can be directly controlled by hyperparameters in Vec2Face. We show that the resulting dataset, HSFace10k, has large intra- and inter-class variations and yields state-of-the-art FR accuracy. Importantly, Vec2Face allows for scaling HSFace to 300K identities and 15M images while still seeing accuracy improvement. In future work, we will study the definition of identity to solve more challenging FR tasks, try other generation structures, and extend this method to generic objects.
@@ A.1 Comparison with Diffusion Model
Training a diffusion model conditioned with image feature vectors is a similar approach to Vec2Face. We evaluate both architectures by measuring the performance on open-set image reconstruction. The Arc2Face (Papantoniou et al., 2024) model, which fine-tuned the stable-diffusion-v1-5 on WebFace42M, is used to represent the diffusion models. The images in an in-the-wild dataset, LFW (Huang et al., 2008), and an indoor dataset, Hadrian (Wu et al., 2024a) are used for evaluation. The features are extracted by ArcFace-R100 and fed to both models. Since both dataset are trained/fine-tuned on the cropped and aligned dataset, no image cropping and alignment are used in this evaluation. Fig. 7 and Fig. 7 show the similarity distribution between the original images and reconstructed images. The observations are: 1) both models can preserve the identity for the in-door images, 2) Vec2Face has better performance on open-set image reconstruction on both datasets, 3) the diffusion model conditioned with image/identity features do not always preserve the identity, especially for in-the-wild images. Hence, the proposed architecture is better than the conventional diffusion model conditioned with image features. Examples are shown in Fig. 19.
@@ A.2 Ablation study of loss functions
A.3 Dataset Noise Analysis
@@
(a) Intra-class Similarity Distributions
(b) Inter-class Similarity Distributions
@@
Following the metric used in previous works (Cao et al., 2018; Wu & Bowyer, 2023; Zhu et al., 2023; Deng et al., 2022), we evaluate the noise using two hard threshold values: one for the similarity between image features and their corresponding identity features, and another for the similarity across identity features. Figure 12(a) shows the cosine similarity distributions of the available synthetic datasets in situation one. Since greater similarity means images are more likely from the same identity, the results show that our dataset has the highest identity consistency within each identity folder. Moreover, generally, if the similarity value of an image feature and its identity feature is less than 0.3, it can be regarded as an outlier/noise (Deng et al., 2022). Hence, the available synthetic datasets contain greater intra-class noise than ours. Figure 12(b) shows the similarity distributions of available synthetic datasets in situation two. WebFace260M (Zhu et al., 2023) is the only work discussing the details of inter-class denoising. It merged two identity folders if their identities have higher than 0.7 similarity, so if the identity similarity is larger than 0.7, we regard it as an inter-class noise. The results show that, except for DCFace and ours, the other datasets have inter-class noise to some extent.
@@
A.4 Age and pose variation
@@
A.5 Feature Interpolation and Feature Value Impact
@@
Figure 17: Impact of feature values analysis. The first row shows images generated with feature values, from left to right, of {0, 1, 1.5, 2, 3, 10, 100}. The second row shows images generated with feature values of {-0.5, -1, -1.5, -2, -3, -10, -100}. The last row shows images generated with feature values of {1, 0.5, 0.25, 0.125, 0.0625, 0.03125, 0.015625}.
@@
We analyze the impact of feature values on the generated images in two ways: 1) interpolating eight feature vectors between two image features and 2) proportionally changing the values in the feature vector. Figure 16 shows the results of feature interpolation between two images. It indicates that the interpolation in the feature domain can be smoothly presented in the image domain, showcasing Vec2Face’s capability of understanding the vector characteristics in the feature domain. Figure 17 shows the results of proportionally changing the values in the feature. The observations are: 1) As the values in a feature vector increase to a large extent, the image stops changing but the quality decreases, for both negative and positive directions; 2) As the values in a feature vector become close to zero, the face attributes are erased and eventually disappear.
A.6 Impact of values at fixed dimensions
@@
Vec2Face draws an image based on the values in the feature vector. We first change the values in one dimension. Fig. 20 shows the cases where some obvious patterns are observed. For each of the four identities, the value at index 15 is related to the hair volume, pose angle, age and expression; the value at index 26 have a high correlation with the bangs, facial hair and age, eyeglasses presence, and expression; increasing the value at index 32 changes the age, head pose, hair color, and expression. The observation is that changing values in a single dimension may change the face attribute but it is not consistent for all identities.
@@
We then change the values in fixed-length (i.e., 8) of dimension chunks and there are some noticeable patterns. Fig. 21 shows the generated images when the weight at specific dimensions increases. A general conclusion is that changing values at dimension chunks varies the facial attributes but the patterns are inconsistent across identities. Specifically, dimension [40:48] changes the age for the first identity, no obvious pattern for the second, the expression for the third, and head pose for the fourth; dimension [56:64] changes the hairstyle for the first identity, facial hair for the second, age and eyeglasses for the third, age and facial hair for the fourth; dimension [96:104] changes the gender for the first identity, age and hairstyle for the second, face exposure level for the third, and facial hair for the fourth. Therefore, it is hard to control the attributes by handcrafting the features. AttrOP is much more efficient and effective.
@@
(a) Changes happen in feature[15]
(b) Changes happen in feature[26]
(c) Changes happen in feature[32]
@@
(a) Changes happen in feature[40:48]
(b) Changes happen in feature[56:64]
(c) Changes happen in feature[96:104]
Figure 21: Examples of changing values in fixed dimensions. From left to right, the value gradually increases in the target dimensions. The raw images are in low quality, so we use AttrOP to increase the image quality.

### v4

@@ 4.4 Further analysis
Vec2Face generates datasets with large inter-class separability. Since only Arc2Face and Vec2Face allow us to generate a large number of identities, we directly use the datasets (including a real dataset) released in previous works to calculate the identity vectors for separability evaluation. For Arc2Face and Vec2Face, we sample 200K well-separated vectors to generate identity images, where we report the results of using LCM-lora (Luo et al., 2023) for Arc2Face. Fig. 5 presents the number of identities whose cosine similarity against any other identity vector is less than 0.4, as more generated identities are gradually added for each dataset. Note that, a reasonable threshold for separability evaluation varies when using different FR models (Kim et al., 2023; Boutros et al., 2024). We select 0.4 is because it gives us a reasonable number (183K out of 200K) on a real dataset. From the results, we have three observations.
Figure 5: Comparing existing synthetic FR dataset generation methods on inter-class separability.
First, synthetic datasets such as SynFace and SFace generally have lower inter-class separability than real-world datasets. This is because their data synthesis methods do not utilize the face feature characteristics for image generation, as Arc2Face and IDiff-Face do. Second, Arc2Face has much larger identity separability than existing methods, comparing with the number reported in (Kim et al., 2023), but is still inferior to Vec2Face. This is because the initial noise image sometimes exerts a stronger influence on the final output than the conditioned identity vector. Third, real-world datasets have slightly lower identity separability than our method, because they have a small number of twins, close relatives, or even the same person with different character names in TV shows (Wu et al., 2024b). Note that 0.4 is not used to evaluate the number of actual identities in the dataset, but rather to provide a consistent measurement of identity separability across datasets. Overall, the special design of Vec2Face help us to generate well-separated identities.
Impact of inter-class separability on FR accuracy. Table 7 summarizes how different average identity similarity or different levels of identity separability affect FR accuracy. Results indicate that high identity separability is beneficial, which is consistent with previous findings (Boutros et al., 2023a). A unique finding is that too large separation does not further benefit the performance.
@@
Table 5: Impact of $\sigma$ values with corresponding sampling % for each identity on average FR accuracy (%). We use 0.5M images from 10K identities for training. AttrOP is not applied and Simmin is the minimum cosine similarity between the perturbed vectors and their ID vectors.
@@
Table 6: Comparing a real dataset with HSFaces on other tasks. Hadrian, Eclipse, SLLFW, and DoppelVer emphasize facial hair variation, face exposure difference, similar-looking, and doppelganger, respectively.
@@
Table 7: Impact of inter-class separability. Larger average identity similarity, or low separability reduces the accuracy. We use 0.5M images from 10K identities for training.
@@
Table 8: Impact of number of images per identity. Different number of images are generated by the best $\sigma$ and their sampling % setting. The average accuracy (%) on five test sets is reported.
Architectures Avg.
@@
Table 9: HSFace10K training on ResNets and Vision Transformers (ViT). The average accuracy (%) on five test sets is reported.
Vec2Face generates datasets with large intra-class variation. The accuracy values in Table 1 indicate the generated dataset has large variation on age and pose. To evaluate the intra-class variation on other attributes, we test the models on Hadrian and Eclipse (Wu et al., 2024a), which have 6K image pairs for each and respectively emphasize variation on facial hair and face exposure. Table 6 shows that the datasets generated by Vec2Face have proper variation on these two attributes, as increasing the dataset size eventually surpasses the accuracy of the real dataset.
Impact of hyperparameter $\sigma$. In Vec2Face, $\sigma$ directly controls intra-class variation, i.e., larger $\sigma$ means higher intra-class variation and vice versa. We test different configurations of $\sigma$ values and its sampling %. Table 5 shows that, when increasing $\sigma$, FR accuracy first increases and then drops. Initially, increasing intra-class variation is beneficial because it creates some hard training samples. But if $\sigma$ is too large, the generated face may look like a different person, degrading the accuracy. Similarly, increasing the sampling % of the large $\sigma$ may hurt the identity consistency in a folder, which also degrades the accuracy. Thus, we choose $\sigma=\{0.3,0.5,0.7\}$ to sample $\{40\%,40\%,20\%\}$ images for each identity.
Does Vec2Face generate new identities compared with its training identities? We compute the feature similarity between the 300K identities from the proposed synthetic data and the 50K training identities selected from WebFace4M. We find that only 0.398%, or 1194 out of the 300K synthetic identities are similar to the training identities, with a similarity score greater than 0.4. To completely avoid using real identities for training FR models, we have replaced these 1194 identities with new identities, and the performance change is negligible. More analysis is in the Appendix A.8.
More images per identity is beneficial in FR accuracy. In Table 8, increasing the number of images per identity from 5 to 80 increases the accuracy, but this improvement is saturated at 80.
Larger backbone helps the FR model performance. In Table 9, accuracy of six models trained with HSFace10K and ArcFace loss increases from SE-IResNet18 to SE-IResNet100 and is saturated at SE-IResNet200. Accuracy increase is also observed when using a larger Vision Transformer (Dosovitskiy, 2020) backbone.
Limitations. The proposed synthetic dataset is effective for general FR but would be less useful for fine-grained tasks such as discriminating between doppelgangers, see Table 6. We believe a better understanding of ‘identity’ and stronger generation models will help solve this issue and we hope to address this in future work.
@@ 5 Conclusions
This paper proposes Vec2Face, a face generation method which can elegantly create a large number of identities and face images. This method conveniently converts vectors to face images and can translate perturbations of the vectors into consistent perturbations of the resulting face images. Because of its unique design, the inter-class and intra-class variations of generated face images can be directly controlled by hyperparameters of vector sampling.. We show that the resulting dataset, HSFace10k, has large intra- and inter-class variations and yields state-of-the-art FR accuracy. Importantly, Vec2Face allows for scaling HSFace to 300K identities and 15M images while still seeing accuracy improvement. In future work, we will study the definition of identity to solve more challenging FR tasks, try other generation structures, and extend this method to generic objects.
@@ A.1 Comparison with diffusion model
Training a diffusion model conditioned with image feature vectors is a similar approach to Vec2Face. Thus, we evaluate both architectures by measuring the performance on open-set image reconstruction. For model choice, Arc2Face (Papantoniou et al., 2024) model, which fine-tunes the stable-diffusion-v1-5 on WebFace42M, is used to represent the diffusion models. For test set choices, the images in an in-the-wild dataset, LFW (Huang et al., 2008), and an indoor dataset, Hadrian (Wu et al., 2024a) are used. Since both model use face feature as input, we extract image features by using a FR model and feed to both models for image reconstruction. Because both dataset are trained/fine-tuned on the cropped and aligned FR datasets, no image cropping and alignment are used in this evaluation. Fig. 7 and Fig. 7 show the similarity distribution between the original images and reconstructed images. The observations are: 1) both models can preserve the identity for the in-door images, 2) Vec2Face has better performance on open-set image reconstruction on both datasets, 3) the diffusion model conditioned with image/identity features do not always preserve the identity, especially for in-the-wild images. Since both model use the features of the same set of images and the variation of the images generated by Arc2Face are only dependent on the initial noise image, we speculate the failed cases of identity preservation are caused by the initial noise image. Hence, the proposed architecture is better than the conventional diffusion model conditioned with image features on identity preservation. Examples are shown in Fig. 16.
@@ A.2 Ablation study of loss functions
A.3 Synthetic FR dataset noise analysis
@@
Following the metric used in previous works (Cao et al., 2018; Wu & Bowyer, 2023; Zhu et al., 2023; Deng et al., 2022), we evaluate the dataset noise using two hard threshold values: one for detecting the outliers within each folder (intra-class noise), and another for detecting similar identities but marked as different in the dataset (inter-class noise). Figure 10 shows the cosine similarity distributions of the available synthetic datasets in situation one. Since greater similarity means images are more likely from the same identity, the results show that our dataset has the highest identity consistency within each identity folder. Moreover, if the similarity value of an image feature and its identity feature is less than 0.3, it can be regarded as an outlier/noise (Deng et al., 2022). With this metric, the available synthetic datasets contain greater intra-class noise than ours. Figure 10 shows the similarity distributions of available synthetic datasets in situation two. WebFace260M (Zhu et al., 2023) is the only work discussing the details of inter-class denoising. It merged two identity folders if their identities have higher than 0.7 similarity, so if the identity similarity is larger than 0.7, we regard it as an inter-class noise. The results show that, except for DCFace and ours, the other datasets have inter-class noise to some extent.
@@
A.4 The effect of age and pose variation of datasets in FR accuracy
@@
A.5 Feature interpolation and feature value impact on generated images
@@
Figure 13: Impact of feature values analysis. The first row shows images generated with feature values, from left to right, of {0, 1, 1.5, 2, 3, 10, 100, 200}. The second row shows images generated with feature values of {-0.5, -1, -1.5, -2, -3, -10, -100, -200}. The last row shows images generated with feature values of {1, 0.5, 0.25, 0.125, 0.0625, 0.03125, 0.015625, 0}.
@@
We analyze the impact of feature values on the generated images in two ways: 1) interpolating six feature vectors between two image features and 2) proportionally changing the values in the feature vector. Figure 12 shows the results of feature interpolation between two images. It indicates that the interpolation in the feature domain can be smoothly presented in the image domain, showcasing Vec2Face’s capability of understanding the vector characteristics in the feature domain. Figure 13 shows the results of proportionally changing the values in the feature. The observations are: 1) As the values in a feature vector increase to a large extent, the image stops changing but the quality decreases, for both negative and positive directions; 2) As the values in a feature vector become close to zero, the face attributes are erased and eventually disappeared.
A.6 Impact of values at dimensions
@@
Vec2Face generates an image based on the values in a feature vector, which motivates an intriguing investigation of how values in dimensions impact the output image. We start at changing the value in one dimension. Fig. 18 shows the cases where some obvious patterns occurring. For each of the four identities, the value at index 15 is related to the hair volume, pose angle, age and expression; the value at index 26 have a high correlation with the bangs, facial hair and age, eyeglasses presence, and expression; increasing the value at index 32 changes the age, head pose, hair color, and expression. The observation is that changing values in a single dimension may change the face attribute but it is not consistent for all identities. In fact, changing value at a single dimension does a negligible change on the generated image in most dimensions.
@@
We then change the values in fixed-length (i.e., 8) of dimension chunks and there are some noticeable patterns. Fig. 17 shows the generated images when the values at specific dimensions increases. A general conclusion is that changing values at dimension chunks varies the facial attributes but the patterns are also inconsistent across identities. For instance, dimension [40:48] changes the age for the first identity, no obvious pattern for the second, the expression for the third, and head pose for the fourth; dimension [56:64] changes the hairstyle for the first identity, facial hair for the second, age and eyeglasses for the third, age and facial hair for the fourth; dimension [96:104] changes the gender for the first identity, age and hairstyle for the second, face exposure level for the third, and facial hair for the fourth. Therefore, it is hard to control the attributes by handcrafting the features. AttrOP is much more efficient and effective.
A.7 Discussion of privacy issue and future dataset usage
Datasets HSFace300K WebFace4M (50K) Glint360K
Accuracy 93.52 96.61 97.63
Table 10: FR accuracy comparison of the model trained with the largest proposed data HSFace300K, the 50K identities used for Vec2Face training, and the pretrained FR model used for feature extraction.
The privacy issue of using images from real identities is the main concern in face recognition technique development. As a result, governments publish regulations (Voigt & Von dem Bussche, 2017) to restrict biometric data usage. Before the possible fully restrictions on real face data usage, although not achieving as good performance as using datasets of real data (see in Table 10), this work provides several datasets for continuing the development of face recognition techniques without violating the regulations. Moreover, the unique design of Vec2Face can be inspired for generating better synthetic datasets in the future.
A.8 Identity leakage
To validate the efficacy of Vec2Face in generating a large number of synthetic identities without causing identity leakage, we conduct two experiments: 1) measuring the identity leakage between identities in the synthetic dataset and WebFace4M identities, and 2) measuring the identity leakage between randomly generated identities and identities in two real datasets. In (Zhu et al., 2023), two identities that have a similarity higher than 0.7 is regarded as the same identity. We use a pre-trained FR model to extract the face feature and adopt the same threshold to measure the identity leakage in percentage. For the former, Table 12 shows that both DCFace and HSFace10K do not have identity leakage issues, but CemiFace has a 24.21% identity leakage. Hence, DCFace and HSFace10K do not violate the regulations, whereas CemiFace does. For the latter, Table 13 indicates that even though we randomly generate 5M identities, Vec2Face does not cause any ID leakage issue, showcasing the potential of effectiveness of synthetic identity generation and dataset scalability.
A.9 Other material
The performance of mixing real and synthetic training sets is in Table 11. The architecture of fMAE is in Figure 15. The FR training configuration is in Table 14. The examples of the effect of stochasticity in fMAE on generated images is in Figure 14.
Real + synthetic Avg.
CASIA-WebFace 94.79
+ HSFace20K 95.67
+ HSFace100K 95.69
+ HSFace200K 95.43
+ HSFace300K 95.72
+ HSFace400K 95.64
Table 11: Average accuracy of the FR model trained with mixed training sets (real + synthetic) tested on LFW, CFP-FP, AgeDB-30, CALFW, and CPLFW.
Dataset ID leakage (%)
DCFace 0%
HSFace10K 0%
CemiFace 24.21%
Table 12: Identity leakage of DCFace, CemiFace (Sun et al., 2024), HSFace10K on WebFace4M dataset. Following the strategy in WebFace4M, we calculate the percentage of synthetic IDs that have a similarity larger than 0.7 with the WebFace4M IDs. CemiFace has a large percent ID leakage.
5M random IDs 0.5 0.7
CASIA-WebFace 0% 0%
WebFace4M 0% 0%
Table 13: ID leakage on 5M randomly sampled IDs. Using the same strategy in Table 12, the results indicate that the Vec2Face model has a huge potential to generate a large number of synthetic identities without causing ID leakage issues.
Figure 14: Examples of the effect of the stochasticity of fMAE. Each vectors is processed 8 times for image generation. The changes on images are negligible.